Thando is 10 years old and his father is 35 years of age.In how many years time will thandos father be twice thando's age?
Answer by Earlsdon(6294)

Thandos = 10
Father = 35
In x year's time, Thandos will be 10+x and his Father will be 35+x, so...
35+x = 2(10+x)
35+x = 20+2x
35 = 20+x
x = 15
In 15 year's time, Thando's Father will be twice Thando's age.
In 15 year's time:
Thandos will be 10+15 = 25 and
his father will be 35+15 = 50